A charge of mass m charge 2e is released from rest in a uniform electric field of strength E .what is the time taken by it to travel a distance d in the field

Grade:12

3 Answers

Adarsh
5 years ago
Dear priyal ,
 
force on charge = 2eE … so acceleration of charge = 2eE/m.
now,  by using  s=ut+1/2at2 
 
we get , d = ½.2eE/m.t2  =>  t = ( md/eE )1/2 
 
Thanka and Regards
Adarsh
Eshan
5 years ago
Dear student,

The force acting on the charge isF=qE=(2e)E

Hence the acceleration of the particle=a=\dfrac{F}{m}=\dfrac{2eE}{m}
From the second equation of motion,

d=ut+\dfrac{1}{2}at^2
\implies d=(0)(t)+\dfrac{1}{2}\dfrac{2eE}{m}t^2

\implies t=\sqrt{\dfrac{dm}{eE}}
Giresh
4 years ago
A Proton of mass M charge e is released from rest in a uniform electric field of strength the time taken by A train travels a distance D in the electric field is √(2dm/eE)